1. Which of the followings are/are not money in the United States? Explain your answers

  1. a US dime
  2. a Mexican peso
  3. a Picasso painting
  4. a plastic credit card
  5. an ounce of gold

2. If the Fed supply 20 billion dollars into the commercial banking system, how much more money the commercial banks create and supply? (rrr = 10%)

3. How would an increase of minimum wage affect unemployment rate?
Answer #1

Answer 1

a US dime is a coin of ten cents hence can be exchanged for goods hence it is a money

b Mexican Peso is a currency of Mexico hence cannot be exchanged for good in US. Hence it is not a money in US

c Picasso Painting cannot be exchanged for good. Hence it is not a money

d A plastic credit card can be used to purchase goods. Hence it is a money

e Ounce of Gold cannot be exchanged for good. Hence it is not a money

Answer 2

Total Money Supply = Initial money Supply/ RRR = 20/10% = 20/0.1 = $200 billion

Hence Banks will create a total money supply of $200 billion

Answer 3

When there is an increase in Minimum Wage, Supply for Labor will increase and Demand for Labor will decrease. Hence the gap between the Supply and demand will increase hence Unemployment will increase.